OpenStudy (anonymous):

finding the slope of an equation of a line -10-x=-2y

OpenStudy (ilovebmth1234):

1/2 i think

OpenStudy (texaschic101):

y = mx + b....the m position is the slope -10 - x = -2y -- divide by -2 5 + 1/2x = y change it if you need to...y = 1/2x + 5 slope = 1/2 .... ilovebmth is correct

OpenStudy (ilovebmth1234):

i did that really quick lol i'm surprised i got that right cause i had to do it in my head :D

OpenStudy (texaschic101):

you did just fine :)